Suppose a random sample of size 40 is selected from a population with

σ = 8.

Find the value of the standard error of the mean in each of the following cases. (Use the finite population correction factor if appropriate. Round your answers to two decimal places.)

(a) The population size is infinite.

(b) The population size is N = 40,000.

(c) The population size is N = 4,000.

(d) The population size is N = 400.
Answer #1

Solution,

Given that,

n = 40

\sigma = 8

a) \sigma \bar x = \sigma / \sqrt n = 8 / \sqrt 40 = 1.26

b) N = 40000

\sigma\bar x = (\sqrt{}N - n / N - 1) *  \sigma / \sqrt n

\sigma\bar x = (\sqrt{}40000 - 40 / 40000 - 1) * 8 / \sqrt 40

\sigma\bar x = 1.26

c) N = 4000

\sigma\bar x = (\sqrt{}N - n / N - 1) *  \sigma / \sqrt n

\sigma\bar x = (\sqrt{}4000 - 40 / 4000 - 1) * 8 / \sqrt 40

\sigma\bar x = 1.25

d) N = 400

\sigma\bar x = (\sqrt{}N - n / N - 1) *  \sigma / \sqrt n

\sigma\bar x = (\sqrt{}400 - 40 / 400 - 1) * 8 / \sqrt 40

\sigma\bar x = 1.20